4. Steves utility function over leisure and consumption is given by NLY) - min (31.7. Wage rate is w and the price of the co intermediate micro
Answer #1

4. a. The consumer's problem is:
max min (32.) s.t. Y+5L = 120
The consumer treats leisure and consumption as perfect complements and consumes them in a fixed ratio: \small Y=3L . Substituting this into the budget equation:
8L = 120 L = 15, Y = 45

b. The consumer's new problem is:
max min (32,Y) s.t. Y + 5L = 120 il L > 15, Y +7.5L = 152.5 if L< 15
Since the consumer chose to work for 8 hours at the original wage rate and this new law would affect the consumer if he were working for more than 8 hours, therefore, this new law does not affect the consumer's choice of leisure and consumption. The consumer will continue to work for 8 hours and consume 45 units.

5. a. The consumer's problem is:
max c s.t. X1 + = 833.33
At equilibrium, marginal rate of substitution is equal to the 'price ratio'.
ܗܘ -cܚ 1236 -1 : 2- assi
Substituting this into the consumer's budget equation:
1230 + 1 = 833.330 833.33 = 4 + 1 = 833.330 + 3 = 1
02 ~ 100.1

b. For the consumer to be a borrower, consumption in period one has to be greater than income in period one:
50O 06

c. The MRS is equal to 'price ratio'.
C1 (1+r) 4/7c2 3/7c1 -1 +rC2= 4
Substituting this into the consumer's budget equation:
9.11r) 400 C1 + 1 = 500++ → C1 = 720+400r. 1+ 180 + 100r
For the consumer to be a borrower, consumption in period one is greater than income in period one:
720 + 400r - 1+ > 500 ) r < 2.2